OpenStudy (anonymous):

CHECK MY ANSWER PLEASE. Solve 3x - 5y = 17 9x - 15y = -4 A. (-1, -4) B. (4, -1) C. No Solutions D. Infinitely Many Solutions<---- I worked out a and b and they don't work so I think it's D

Directrix (directrix):

3x - 5y = 17 --Multiply by 3 --> 9x - 15y = 51 9x - 15y = -4 Note that the second equation has the same left side as the first. The constant terms on the right are different. What does that mean?

OpenStudy (anonymous):

That there are no solutions?

Directrix (directrix):

Correct. If you graph the lines, you will see that they are parallel and do not intersect. So, no solution.

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

Thank you!! Could you help me with a few more? I don't know how to do one with a fraction

Directrix (directrix):

Close this thread and open a new one with the new question. I'll try to help.